# Language model developers should report train-test overlap

### v1

@@ Abstract
Language models are extensively evaluated, but correctly interpreting evaluation results requires knowledge of train-test overlap, which refers to the extent to which the language model is trained on the very data it is being tested on. The public currently lacks adequate information about train-test overlap: most models have no public train-test overlap statistics, and third parties cannot directly measure train-test overlap since they do not have access to the training data. To make this clear, we document the practices of 30 models, finding that just 9 models report train-test overlap: 4 models release training data under open-source licenses, enabling the community to directly measure train-test overlap, and 5 models publish their train-test overlap methodology and statistics. By engaging with language model developers, we provide novel information about train-test overlap for three additional models. Overall, we take the position that language model developers should publish train-test overlap statistics and/or training data whenever they report evaluation results on public test sets. We hope our work increases transparency into train-test overlap to increase the community-wide trust in model evaluations.
@@ 1 Introduction
A growing literature demonstrates high train-test overlap for language models, which contributes to significant degradation in performance between seen and unseen test examples (Lewis et al., 2020; Elangovan et al., 2021; Vu et al., 2023). For example, OpenAI initially reported that GPT-4 had achieved state-of-the-art performance on a test set of Codeforces coding questions, claiming that there was no contamination (OpenAI, 2023). Yet it was later demonstrated [^1] See https://twitter.com/cHHillee/status/1635790330854526981?lang=en that while GPT-4 achieves 100% accuracy for 10 pre-2021 problems, the model achieves 0% accuracy on more recent problems. More generally, Kapoor & Narayanan (2022) document that test data often leaks into training data across many domains. Therefore, improving the community’s understanding train-test overlap will increase the validity of, and trust in, evaluations.
Given the value of understanding train-test overlap, we study the practices of 30 language model. We find that 9 models have published sufficient data for the AI community to contextualize train-test overlap: 4 models (OLMo—AI2, GPT-NeoX—EleutherAI, RedPajama INCITE—Together, StarCoder 2—BigCode/HuggingFace/ServiceNow) have released open-source datasets that the community can inspect for train-test overlap, and 5 models (GPT-4—OpenAI, Llama 3.1—Meta, Qwen2—Alibaba, Palmyra—Writer, Apple Intelligence—Apple) have published their methodology and statistics for train-test overlap. The remaining 23 models do report evaluation results on public test sets, but do not (adequately) report train-test overlap results. In parallel to models reporting train-test overlap, the community is building black-box methods to estimate train-test overlap without access to training data (Golchin & Surdeanu, 2023; Shi et al., 2023; Oren et al., 2023), but these approaches are quite limited at present. We take the position that language model developers should report train-test overlap.
@@
2 Strategies to Estimate and Address Train-Test Overlap
@@
involve researchers working to estimate train-test overlap through model api access and the test set rather than directly through access to the training set. Notably, there have been efforts to estimate train-test overlap via prompting, word probabilities, and test example orderings (Golchin & Surdeanu, 2023; Shi et al., 2023; Oren et al., 2023). Golchin & Surdeanu (2023) prompt the model with the dataset name, partition type, and an initial segment of the reference string, and mark train-test overlap if the model responds with a exact or similar copy in the output. Shi et al. (2023) estimate train-test overlap via the probability outputs of outlier words, with the hypothesis that unseen examples is likely to contain few outliers with low probabilities. Oren et al. (2023) estimate train-test overlap by considering the ordering of test instances, noting that language models with train-test overlap are likely to memorize such ordering. These methods can be helpful for estimation and as a sanity check to white-box approaches, but have currently have limitations as they are not robust to adversarial settings such as if a developer fine-tuned its model to avoid revealing training data and even in the benign setting, require certain assumptions such as requiring a certain threshold of frequencies for detection or certain methods of training (Casper et al., 2024; Golchin & Surdeanu, 2023; Shi et al., 2023; Oren et al., 2023). Estimating and interpreting train-test overlap is difficult even in the white-box setting with direct access to the training data as current approaches have significant limitations; with further constraints in the black-box setting, the challenges only increase.
@@
that include data that was produced after the knowledge cutoff date of a model also help mitigate train-test overlap. Including recent data is a best practice for new test sets, though this may be difficult if, for example, a new test set is derived from existing data (e.g. based on old Wikipedia data or AI-generated data). Even when this approach is implemented successfully, new models are released regularly that are trained on more recent data, necessitating some analysis of train-test overlap with the previously novel test set. One modification of this approach is to add novel data to the test set at regular intervals, as with Livebench (White et al., 2024) or Image2Struct. [^3] See https://crfm.stanford.edu/helm/image2struct/latest/ In addition to the financial cost of continually adding novel data, which may not be feasible for every domain or project, one challenge of this approach is that it is difficult to interpret longitudinal progress.
@@ 3 Language Models
To establish a broad understanding of the landscape, we comprehensively consider the train-test overlap practices of the flagship language model of 30 developers (01.ai, Adept, AI2, AI21 Labs, Aleph Alpha, Alibaba, Amazon, Anthropic, Apple, BigCode, Cohere, Databricks, DeepSeek, EleutherAI, Technology Innovation Institute, Google, IBM, Imbue, Inflection, Meta, Microsoft, Mistral, NVIDIA, OpenAI, Reka AI, Snowflake, Stability AI, Together AI, Writer, and xAI). We assembled this list by considering all models on the HELM MMLU leaderboard [^4] See https://crfm.stanford.edu/helm/mmlu/latest/#/leaderboard and additional models that we selected for impact and relevance based on Ecosystem Graphs (Bommasani et al., 2023b).
@@ 4.3 Scores
Here we document the train-test overlap practices of 30 models with published results on public test sets. Of these, 9 models have published sufficient data for the community to contextualize train-test overlap: 4 models (OLMo—AI2, GPT-NeoX—EleutherAI, RedPajama INCITE—Together, StarCoder 2—BigCode/HuggingFace/ServiceNow) models have released training data under open-source licenses, which researchers can inspect for train-test overlap and 5 models (GPT-4—OpenAI, Llama 3.1—Meta, Qwen2—Alibaba, Palmyra—Writer, Apple Intelligence—Apple) have published their methodology and statistics for train-test overlap. For developers that do not openly release their training data, we provide additional explanation below as to why their transparency regarding train-test overlap is meaningful; for quantification of the degree of train-test overlap for each of these developers, see their associated technical reports.
OpenAI
Meta
Writer
released train-test overlap statistics for Palmyra X after receiving a request from the authors. Writer ran a script [^6] This script is publicly available at https://github.com/stanford-crfm/data-overlap—we encourage other developers to run it over their training sets and publicly report the results. provided by the authors over its pretraining and instruction-tuning data to evaluate train-test overlap via an n-gram analysis; the results of this analysis are accessible at https://drive.google.com/file/d/1-3_GtZTyIbE1B5XWGGWtLRy5OwVtG0EK/view?usp=sharing. Writer publishes its results on HELM Lite (Writer, 2024), which includes 9 public test sets, and Writer reported train-test overlap on each of the public test sets included in HELM as well as others. Writer found some degree of train-test overlap for 13 of the 27 test sets on which it ran the script (APPS, CivilComments, CNN/Daily Mail, EntityMatching, HumanEval, ICE, LegalSupport, MATH, NarrativeQA, RAFT, The Pile, WikiFact, XSum).
Alibaba
Apple
@@ 5 Discussion
Overall, while train-test overlap is a fundamental to interpreting evaluation results, there is still significant limitations in the measurement methodology, beyond data access challenges and developer responsibility. As described above, direct string comparison is the most common way to quantify train-test overlap. This method has slight variations, but typically involves detecting substring matches between training and test data. N-gram matching is commonly employed (Yang et al., 2024; Dubey et al., 2024; Brown et al., 2020; Chowdhery et al., 2022), where documents are tokenized and then compared, though OpenAI compares characters rather than tokens in its analysis for GPT-4 (OpenAI, 2023). There are important design decisions developers make in employing n-gram strategies, such as what to set as N, whether to allow fuzzy matching or skipgrams (Dubey et al., 2024), and whether to filter based only a threshold of matches. This lack of uniformity in measurement approaches can make it challenging to directly compare train-test overlap analyses.
@@
There are a number of limitations to this class of approaches. One limitation is that n-grams are coarse and do not capture the differences in types of overlap between different test sets. For instance, CivilComments is derived from news sites, so overlap between training data and CivilComments is likely due to news articles that appear in both the training and test data (Duchene et al., 2023). In contrast, MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b) are in question-answer format, so overlap could stem from leakage of questions or answers or repetition of common phrases in questions and answers. We categorize the overlap types for different scenarios for The Pile (Gao et al., 2020) in 2. Overlap types can be broadly categorized into question leakage; quotes from news, laws, books, and songs; common phrases; and multi-token identifiers. Question leakage is the canonical concern for training data: if a model trains on the questions in the test set, it can achieve high results that fail to generalize well to new questions. However, these other overlap types can also be informative; for instance, simply matching the LSAT question stem “Which one of the following could be a complete and accurate list of the” suggests that the training data likely contains LSAT questions or similar questions. Indeed, train-test overlap is a construct that captures the relation between train and test data, and the different type of overlap add complexity that make it difficult to capture in a single statistic. Future work could explore these various overlap types in more detail and devise more granular metrics of measurement.
@@
Additionally, there are complexities in determining what qualifies as the training set, as there are often multiple stages of training and datasets, including pretraining, fine-tuning, and safety alignment among others (Yang et al., 2024; Dubey et al., 2024; Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023). This is often not captured in developers’ public reports (Bommasani et al., 2024a), and it may not be well captured internally either. Precision about the training set is important, though it is beyond the scope of this paper.
This paper does not assert a position on which method is best, and acknowledges that there is substantial research remaining to investigate better methods of computing train-test overlap. Nevertheless, the limitations of black-box approaches are far greater than those of white-box approaches. Just as the Foundation Model Transparency Index has helped improve the transparency of foundation model developers (Bommasani et al., 2023a; 2024a), our hope is that an increase in the number of model developers that report train-test overlap will produce better methods of measurement and help standardize reporting such that developers’ transparency on train-test overlap improves.

### v2

@@ Abstract
Language models are extensively evaluated, but correctly interpreting evaluation results requires knowledge of train-test overlap, which refers to the extent to which the language model is trained on the very data it is being tested on. The public currently lacks adequate information about train-test overlap: most models have no public train-test overlap statistics, and third parties cannot directly measure train-test overlap since they do not have access to the training data. To make this clear, we document the practices of 30 models, finding that just 9 models report train-test overlap: 4 models release training data under open-source licenses, enabling the community to directly measure train-test overlap, and 5 models publish their train-test overlap methodology and statistics. By engaging with language model developers, we provide novel information about train-test overlap for three additional models. Overall, this position paper argues that language model developers should publish train-test overlap statistics and/or training data whenever they report evaluation results on public test sets. We hope our work increases transparency into train-test overlap to increase the community-wide trust in model evaluations.
@@ 1 Introduction
A growing literature demonstrates high train-test overlap for language models, which contributes to significant degradation in performance between seen and unseen test examples (Lewis et al., 2020; Elangovan et al., 2021; Vu et al., 2023). For example, OpenAI initially reported that GPT-4 had achieved state-of-the-art performance on a test set of Codeforces coding questions, claiming that there was no contamination (OpenAI, 2023). Yet it was later demonstrated [^1] See https://twitter.com/cHHillee/status/1635790330854526981?lang=en that while GPT-4 achieves 100% accuracy for 10 pre-2021 problems, the model achieves 0% accuracy on more recent problems. More recently, Anthropic claimed a "zero to one" moment in solving Capture The Flag tasks (Anthropic, 2025). Yet Transluce found that Claude 3.5 solves broken tasks by memorizing answers and hallucinating them (Meng, 2025). More generally, Kapoor & Narayanan (2023) document that test data often leaks into training data across many domains. Therefore, improving the community’s understanding of train-test overlap will increase the validity of, and trust in, evaluations.
Given the value of understanding train-test overlap, we study the practices of 30 language models. We find that 9 models have published sufficient data for the AI community to contextualize train-test overlap: 4 models (OLMo—AI2, GPT-NeoX—EleutherAI, RedPajama INCITE—Together, StarCoder 2—BigCode/HuggingFace/ServiceNow) have released open-source datasets that the community can inspect for train-test overlap, and 5 models (GPT-4—OpenAI, Llama 3.1—Meta, Qwen2—Alibaba, Palmyra—Writer, Apple Intelligence—Apple) have published their methodology and statistics for train-test overlap. The remaining 23 models do report evaluation results on public test sets, but do not (adequately) report train-test overlap results. In parallel to models reporting train-test overlap, the community is building black-box methods to estimate train-test overlap without access to training data (Golchin & Surdeanu, 2023; Shi et al., 2023; Oren et al., 2023), but these approaches are quite limited at present. We take the position that language model developers should report train-test overlap.
@@
2 Alternative Views
@@
involve researchers working to estimate train-test overlap through model API access and the test set rather than directly through access to the training set. Notably, there have been efforts to estimate train-test overlap via prompting, word probabilities, and test example orderings (Golchin & Surdeanu, 2023; Shi et al., 2023; Oren et al., 2023). Golchin & Surdeanu (2023) prompt the model with the dataset name, partition type, and an initial segment of the reference string, and mark train-test overlap if the model responds with an exact or similar copy in the output. Shi et al. (2023) estimate train-test overlap via the probability outputs of outlier words, with the hypothesis that unseen examples is likely to contain few outliers with low probabilities. Oren et al. (2023) estimate train-test overlap by considering the ordering of test instances, noting that language models with train-test overlap are likely to memorize such ordering. These methods can be helpful for estimation and as a sanity check to white-box approaches, but currently have limitations as they are not robust to adversarial settings such as if a developer fine-tuned its model to avoid revealing training data and even in the benign setting, require certain assumptions such as requiring a certain threshold of frequencies for detection or certain methods of training (Casper et al., 2024; Golchin & Surdeanu, 2023; Shi et al., 2023; Oren et al., 2023). Estimating and interpreting train-test overlap is difficult even in the white-box setting with direct access to the training data as current approaches have significant limitations; with further constraints in the black-box setting, the challenges only increase.
@@
that include data that was produced after the knowledge cutoff date of a model also help mitigate train-test overlap. Including recent data is a best practice for new test sets, though this may be difficult if, for example, a new test set is derived from existing data (e.g. based on old Wikipedia data or AI-generated data). Even when this approach is implemented successfully, new models are released regularly that are trained on more recent data, necessitating some analysis of train-test overlap with the previously novel test set. One modification of this approach is to add novel data to the test set at regular intervals, as with Livebench (White et al., 2024) or Image2Struct(Roberts et al., 2024). In addition to the financial cost of continually adding novel data, which may not be feasible for every domain or project, one challenge of this approach is that it is difficult to interpret longitudinal progress.
@@
3 Consequences of Not Reporting Train-Test Overlap
Failing to disclose train-test overlap can degrade trust within both the AI community and the broader public, ultimately diminishing the value of evaluations that are pivotal for research progress and real-world applications. The current landscape of AI evaluation is increasingly characterized by accusations of "cheating" or undisclosed advantages, fostering a climate of low trust within the community . This lack of transparency not only undermines the credibility of individual model evaluations but also has broader negative consequences for the advancement and trustworthiness of AI as a whole.
One significant consequence of not reporting train-test overlap is that it leads to unsubstantiated claims by model developers, and consequently an erosion of trust. For example, OpenAI initially reported that GPT-4 had achieved state-of-the-art performance on a test set of Codeforces coding questions, claiming that there was no contamination (OpenAI, 2023). Yet it was later demonstrated [^3] See https://twitter.com/cHHillee/status/1635790330854526981?lang=en that while GPT-4 achieves 100% accuracy for 10 pre-2021 problems, the model achieves 0% accuracy on more recent problems. More recently, Anthropic claimed a "zero to one" moment in solving Capture The Flag tasks (Anthropic, 2025). Yet Transluce found that Claude 3.5 solves broken tasks by memorizing answers and hallucinating them (Meng, 2025). Instead, if the model developers had provided train-test overlap statistics associated with these results, it would have been clear that train-test overlap was a significant contributor of these breakthrough results. Accordingly, it increases accountability of model developer, which then increases the trust in their claims.
Perhaps even more significant than these instances with clear evidence are accusations of cheating that are hard to verify, which end up eroding trust further. For instance, there are "cheating" accusations against o3 on the ARC-AGI benchmark (Lee, 2024), which OpenAI has denied. While we can attempt to derive evidence via black-box methods (Golchin & Surdeanu, 2023; Shi et al., 2023; Oren et al., 2023), they have limitations and cannot fully substitute for transparent reporting from model developers. The inherent uncertainty associated with black-box methods, coupled with the potential for adversarial manipulation, means that accusations and counter-accusations can persist, leading to a fragmented and less productive research environment.
Accordingly, the failure to report train-test overlap has significant negative consequences for the AI community. It fosters a climate of distrust, hinders scientific progress by obscuring the true nature of model capabilities, and makes it difficult to build upon existing research. By embracing transparency and routinely reporting train-test overlap statistics, the AI community can move towards a more trustworthy and collaborative environment that ultimately accelerates the development of safer and more trustworthy AI.
@@ 4 Language Models
To establish a broad understanding of the landscape, we comprehensively consider the train-test overlap practices of the flagship language model of 30 developers (01.ai, Adept, AI2, AI21 Labs, Aleph Alpha, Alibaba, Amazon, Anthropic, Apple, BigCode, Cohere, Databricks, DeepSeek, EleutherAI, Technology Innovation Institute, Google, IBM, Imbue, Inflection, Meta, Microsoft, Mistral, NVIDIA, OpenAI, Reka AI, Snowflake, Stability AI, Together AI, Writer, and xAI). We assembled this list by considering all models on the HELM MMLU leaderboard [^4] See https://crfm.stanford.edu/helm/mmlu/v1.8.0/ on October 7, 2024. and additional models that we selected for impact and relevance based on Ecosystem Graphs (Bommasani et al., 2023b).
@@ 5.3 Scores
Here we document the train-test overlap practices of 30 models with published results on public test sets. Of these, 9 models have published sufficient data for the community to contextualize train-test overlap: 4 models (OLMo—AI2, GPT-NeoX—EleutherAI, RedPajama INCITE—Together, StarCoder 2—BigCode/HuggingFace/ServiceNow) models have released training data under open-source licenses, which researchers can inspect for train-test overlap and 5 models (GPT-4—OpenAI, Llama 3.1—Meta, Qwen2—Alibaba, Palmyra—Writer, Apple Intelligence—Apple) have published their methodology and statistics for train-test overlap. For model developers that do not openly release their training data, see Appendix A for additional explanation below as to why their transparency regarding train-test overlap is meaningful.
@@ 6 Discussion
Overall, while train-test overlap is a fundamental to interpreting evaluation results, there is still significant limitations in the measurement methodology, beyond data access challenges and developer responsibility. As described above, direct string comparison is the most common way to quantify train-test overlap. This method has slight variations, but typically involves detecting substring matches between training and test data. N-gram matching is commonly employed (Yang et al., 2024; Dubey et al., 2024; Brown et al., 2020b; Chowdhery et al., 2022a), where documents are tokenized and then compared, though OpenAI compares characters rather than tokens in its analysis for GPT-4 (OpenAI, 2023). There are important design decisions developers make in employing n-gram strategies, such as what to set as N, whether to allow fuzzy matching or skipgrams (Dubey et al., 2024), and whether to filter based only a threshold of matches. This lack of uniformity in measurement approaches can make it challenging to directly compare train-test overlap analyses.
@@
There are a number of limitations to this class of approaches. One limitation is that n-grams are coarse and do not capture the differences in types of overlap between different test sets. For instance, CivilComments is derived from news sites, so overlap between training data and CivilComments is likely due to news articles that appear in both the training and test data (Duchene et al., 2023). In contrast, MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b) are in question-answer format, so overlap could stem from leakage of questions or answers or repetition of common phrases in questions and answers. We categorize the overlap types for different scenarios for The Pile (Gao et al., 2020) in Table 2. Overlap types can be broadly categorized into question leakage; quotes from news, laws, books, and songs; common phrases; and multi-token identifiers. Question leakage is the canonical concern for training data: if a model trains on the questions in the test set, it can achieve high results that fail to generalize well to new questions. However, these other overlap types can also be informative; for instance, simply matching the LSAT question stem “Which one of the following could be a complete and accurate list of the” suggests that the training data likely contains LSAT questions or similar questions. Indeed, train-test overlap is a construct that captures the relation between train and test data, and the different type of overlap add complexity that make it difficult to capture in a single statistic. Future work could explore these various overlap types in more detail and devise more granular metrics of measurement.
@@
Additionally, there are complexities in determining what qualifies as the training set, as there are often multiple stages of training and datasets, including pretraining, fine-tuning, and safety alignment among others (Yang et al., 2024; Dubey et al., 2024; Brown et al., 2020b; Chowdhery et al., 2022a; OpenAI, 2023). This is often not captured in developers’ public reports (Bommasani et al., 2024a), and it may not be well captured internally either. Precision about the training set is important, though it is beyond the scope of this paper.
This paper does not assert a position on which method is best, and acknowledges that there is substantial research remaining to investigate better methods of computing train-test overlap. Nevertheless, the limitations of black-box approaches are far greater than those of white-box approaches. Just as the Foundation Model Transparency Index has helped improve the transparency of foundation model developers (Bommasani et al., 2023a, 2024a), our hope is that an increase in the number of model developers that report train-test overlap will produce better methods of measurement and help standardize reporting such that developers’ transparency on train-test overlap improves.
7 Future Work: Standardization of Train-Test Overlap
The first critical step toward addressing train-test overlap is establishing a baseline expectation for model developers to publish overlap statistics for the benchmarks they themselves report, which we have advocated for here. We have seen promising adoption of this by nine model developers, including three new model developers from outreach. As this expectation becomes increasingly a reality, we have substantial additional work to further increase train-test overlap transparency.
In particular, the model developers employ different methodology for computing train-test overlap statistics, which make it challenging to compare statistics between models. In fact, train-test overlap statistics can differ even between models for a given model developer, without clear reasoning. As such, we believe that standardization of the protocol to compute train-test overlap will be an essential next step to improving transparency and trust in the space.
Here, we document an initial effort where we advocated for a standardized protocol requiring developers to compute overlap using shared benchmarks, tools, and scripts to ensure comparability (see Appendix B-D for more details). While this approach garnered interest among a few model developers including Writer, Meta, and Cohere, widespread adoption proved challenging due to a lack of incentives to promote transparency.
Our hope is that the community can learn from our efforts in standardization to better improve transparency in this space. Our view is that as community awareness of and discussion of train-test overlap becomes increasingly prevalent, model developers will have stronger incentives to release train-test overlap statistics and work toward standardization. With our efforts, we have taken initial steps in this direction, but there is still a significant amount of work to be done.
@@ 8 Conclusion
Impact Statement
Transparency of train-test overlap is essential for community-wide trust in model evaluation. In this work, we seek to increase transparency by urging language model developers to publish train-test overlap statistics and/or training data whenever they report evaluation results on public test sets.
Appendix A Scoring Details
For developers that do not openly release their training data, we provide additional explanation below as to why their transparency regarding train-test overlap is meaningful; for quantification of the degree of train-test overlap for each of these developers, see their associated technical reports.
OpenAI
Meta
Writer
released train-test overlap statistics for Palmyra X after receiving a request from the authors. Writer ran a script  [^6] This script is publicly available and attached in the supplementary material–we encourage other developers to run it over their training sets and publicly report the results. provided by the authors over its pretraining and instruction-tuning data to evaluate train-test overlap via an n-gram analysis. Writer publishes its results on HELM Lite (Writer, 2024), which includes 9 public test sets, and Writer reported train-test overlap on each of the public test sets included in HELM as well as others. Writer found some degree of train-test overlap for 13 of the 27 test sets on which it ran the script (APPS, CivilComments, CNN/Daily Mail, EntityMatching, HumanEval, ICE, LegalSupport, MATH, NarrativeQA, RAFT, The Pile, WikiFact, XSum).
Alibaba
Apple
Appendix B Protocol
Figure 1: Protocol for computing train-test overlap. The first party takes private training data and public test data to compute overlap statistics at the instance level. This is sent to the third party, which aggregates statistics to the dataset level and publishes the results publicly.
As the training data for many language models is private, we propose a protocol (see Figure 1) to coordinate between the model provider (first party) and the external entity (third party) to report train-test overlap statistics. Through this protocol, relevant train-test overlap statistics can be made public while respecting the access controls placed on the underlying training data.
B.1 Notation
The training data, denoted $D_{\text{train}}$, is a set of examples where each $x_{\text{train}}\in D_{\text{train}}$ is a sequence of tokens. The test data consists of one or more test sets $D_{\text{test}}$, where $D_{\text{test}}$ is a set of examples where each instance $(x_{\text{test}},Y_{\text{test}})\in D_{\text{test}}$ consists of an input sequence of tokens $x_{\text{test}}$ and a reference set of token sequences $y_{\text{test}}\in Y_{\text{test}}$ representing the ground truth response. For instance, for a Q&A test set, input would be the question whereas reference are the answer choices. We concatenate the reference set into a single sequence, so we will refer to the reference set as a single reference token sequence $y_{\text{test}}$ hereafter.
B.2 Steps
We divide the above described process of computing train-test overlap metrics between a first-party actor with access to the training set and a third-party actor in the following manner (depicted in Figure 1):
  1. 1.
    The third-party actor(s) release test data, compute-instance-overlap-statistics, aggregate-overlap-statistics, aggregate-metrics publicly to instantiate the protocol.
  2. 2.
    The first-party actor uses compute-instance-overlap-statistics on their $D_{\text{train}}$ and publicly available test data to output instance-level overlap statistics.
  3. 3.
    The third-party actor uses aggregate-overlap-statistics on the instance-level overlap statistics to produce test-set-level statistics, which are published as meaningful overlap information between $D_{\text{train}}$ and test data.
In terms of costs, we distinguish compute-instance-overlap-statistics as the computationally expensive step relative to aggregate-overlap-statistics as training sets are often several terabytes of data. This enables rapid iteration with different aggregation methods on the instance-level overlap stats.
Appendix C Protocol Instantiation
While this protocol is generic with respect to the particular method of computing train-test overlap, we instantiate it with n-gram overlap as it is the predominant method used by model providers and computationally inexpensive (Brown et al., 2020a; Touvron et al., 2023). Intuitively, for large enough $n$, the higher the overlap between the n-grams in $D_{\text{train}}$ and the n-grams in $D_{\text{test}}$ the larger the likelihood that the training set is “contaminated", i.e., contains significant portions of the test set. A variety of metrics aimed at quantifying this intuition were proposed in earlier works, each proposing a different metric over the intersection between n-grams from $D_{\text{train}}$ and $D_{\text{test}}$. We denote the family of such overlap metrics as $\mathcal{F}$, and conceptually organize some of the previously proposed overlap metrics $f\in\mathcal{F}$ in Section LABEL:sec:metrics. Below, we use the abstract form $f\in\mathcal{F}$ and instantiate the protocol in Figure 1 using n-gram overlap with the following test data, compute-instance-overlap-statistics, aggregate-overlap-statistics:
  1. 1.
    The instantiated test data are HELM scenarios (Liang et al., 2022).
  2. 2.
[…]
  3. 3.
    The instantiated aggregate-overlap-statistics takes in the per-instance scores and then aggregates the statistics for public release. We present these aggregations in the results section. These aggregations are the only public information released regarding $D_{\text{train}}$.
Note that there are additional complexities in terms of the protocol. For instance, different providers are comfortable with sharing varying levels of instance information. For simplicity, we note that almost all providers were willing to share instance-level metrics without sharing instance-ids or n-grams. One provider was willing to directly share n-grams, which enabled us to compute metrics on our own without their input (hence saving their time and reducing the need for us to reach out after modifying the protocol). With another, we were only able to run an earlier version of the protocol where we had only binary overlap and have not been able to successfully rerun the protocol since.
Appendix D Train-test overlap metrics
In this section we formally present the train-test overlap metrics. We begin by surveying the existing methods that were used for computing $n$-gram based overlap metrics for leading LLMs, and from which we derive three core metrics.
D.1 Existing Methods
Table 3: Overlap Metrics of Existing Models. Here we note how train-test overlap was computed for selected existing models. In addition to analyzing overlap after training, certain providers filtered training or test data based on overlap. All methods are n-gram based using tokens, besides GPT-4 which is based on characters.
Model Measurement Type Stage
Llama 2 (Touvron et al., 2023) $>$10 tokens Token Post-training analysis
PALM (Chowdhery et al., 2022b) 8 tokens Jaccard Post-training analysis
GPT-3 (Brown et al., 2020a) 13 tokens Binary Pre-training filtering
GPT-3 (Brown et al., 2020a) 8-13 tokens Binary Post-training analysis
GPT-4 (OpenAI, 2023) 50 characters Binary Post-training analysis
Gopher (Rae et al., 2022) 13 tokens Jaccard Pre-training filtering
OPT-IML (Iyer et al., 2023) 13 tokens Binary Post-training filtering
Megatron GPT2 (Shoeybi et al., 2020) 8 tokens Binary Post-training analysis
We surveyed literature on how providers have evaluated train-test overlap on existing LMs to establish a baseline for what providers are familiar and comfortable with. Table 3 contains information on how leading LLMs have computed overlap. All the methods were ngram-based and 13-tokens was the most common. From n-grams, providers computed either binary, jaccard, or token overlap. Binary overlap is the most common and simplest method, which simply marks an instance as overlapping if there is a single overlapping n-gram, and not overlapping otherwise. In contrast, jaccard and token overlap compute a score based on what portion of a given instance is overlapping. Jaccard overlap takes the fraction of the number of n-grams that are overlapping over the total number of n-grams for a given test example. Token overlap counts the number of overlapping tokens over the total number of tokens for a given test instance, which avoids double counting any given token. We provide an example of binary, jaccard, and token overlap below and then define the metrics more formally later. We compute binary, jaccard, and token overlap for the training data as they can be derived from the same primitives (n-grams) and thus require minimal additional compute.
Note that there are additional variations that we do not capture here. For instance, the GPT-3 prefiltering stage computed the frequency of 13-grams within the training set and filtered out those with a frequency greater than 10 (Brown et al., 2020a). We explored filtering and weighting by frequency, but did not find the effect to be sufficiently meaningful to justify the added complexity. We do not capture other variants, e.g. GPT-4 only subsamples 3 50-character segments (OpenAI, 2023) and Llama 2 introduces skipgrams (Touvron et al., 2023)
{mdframed}
Example of Overlap Metrics:
Example sentence: “this is a fake example sentence for showing how we compute metrics”
Example overlapping n-grams (for $n=3$): [(“this is a”), (“is a fake”), (“for showing how”)]   Metric Value Explanation Binary overlap 1 There is at least one overlapping n-gram. Jaccard overlap $\nicefrac{{3}}{{10}}$ There are 3 overlapping n-grams out of 10 total n-grams. Token overlap $\nicefrac{{7}}{{12}}$ There are 7 overlapping tokens out of a total of 12 tokens.
D.2 Metric definition
Based on our literature review, we choose to compute binary, jaccard, and token overlap on 13-grams.
We define the above metrics formally by specifying various overlap metric functions $f\in\mathcal{F}$, where $\mathcal{F}$ is the family of functions $f$ that take a set of sequences of tokens $D_{1}$ and a token sequence $x$ as input and produce a real value between 0 and 1.
We take $D_{1}$ as $D_{\text{test}}$ and $x\in D_{\text{train}}$. When running the algorithm, we load the n-grams associated with the entire $D_{\text{test}}$ into memory and iterate through the test set n-gram by n-gram to compute overlap. While we’ve previously noted that instances in $D_{\text{test}}$ can have two components, input and reference, we will define metrics in terms of token sequence $x$ for simplicity.
Binary overlap marks an instance as overlapping if there is at least a single overlapping n-gram. Mathematically, it is defined as:
$f_{\text{binary}}(D_{\text{train}},x)=\min(|\text{n-grams}(x)\cap\text{n-grams}(D_{\text{train}})|,1)$
Jaccard overlap measures how many n-grams are overlapping for an instance and divides by the total number of n-grams in that instance. Jaccard overlap is defined as:
$f_{\text{Jaccard}}(D_{\text{train}},x)=\frac{|\text{n-grams}(x)\cap\text{n-grams}(D_{\text{train}})|}{|\text{n-grams}(x)|}$
Token overlap measures how many tokens are overlapping for an instance and divides by the total number of tokens in that instance. A token is overlapping if it is associated with at least one overlapping n-gram. This is similar to jaccard, but does not double count tokens for contiguous n-grams. Mathematically: Let $\text{tokens}(x)$ denote the tokens $x_{1},x_{2},...,x_{k}$ corresponding to $x$.
[…]
$f_{\text{Token}}(D_{\text{train}},x)=\frac{|\text{tokens-intersect}(x,D_{\text{train}})|}{|\text{tokens}(x)|}$
D.3 Metric Aggregation
We aggregate the metrics in two ways. Possible overlap instances are those with any overlap in either the input or reference, i.e., binary overlap value of 1. We report the scores separately between input and reference rather than the union. Likely overlap instances are those where both the input and reference overlap of an instance is "dirty". Similar to Touvron et al. (2023), we define "dirty" as a token overlap score of 0.8 or greater, though we simply define any other score as "clean" for simplicity.
D.4 Algorithm
Algorithm for computing overlapping n-grams and frequencies. Code will be released on GitHub.
Algorithm 1 Compute Overlapping N-grams
0:  Test set examples $x_{\text{test}}\in D_{\text{test}}$
0:  Train set examples $x_{\text{train}}\in D_{\text{train}}$
1:  Read the test set examples $x_{\text{train}}\in D_{\text{train}}$ into a hash table $h$ into memory
2:  for $x_{\text{train}}\in D_{\text{train}}$ do
3:     Split it into n-grams $n_{1},\ldots,n_{k}$
4:     for each n-gram $n_{j}$ do
5:        if $n_{j}$ exists in the hash table $h$ then
6:           Mark $n_{j}$ as overlapping
7:        end if
8:     end for
9:  end for
10:  Output the overlapping n-grams, and associated test examples